# Recurrent Complex-Weighted Autoencoders for Unsupervised Object Discovery

### v2

@@ 5 Conclusion
We explored synchrony-based binding with an architecture, SynCx, that differs from current models in three respects: i) SynCx’s weights are complex valued, allowing it to encode joint feature-phase configurations in the weights; ii) SynCx is recurrent and stateful, allowing it to perform iterative constraint propagation; and iii) SynCx has an internal representational bottleneck, which require it to use the phases of complex-valued activations to encode statistical regularities in images. These three properties are crucial to extending object-centric phase synchronization behavior to the fully unsupervised setting. Our conceptually elegant model outperforms all state-of-the-art baselines on Tetrominoes and overcomes a common failure mode of these baselines in which similarly colored objects are grouped together. Overreliance on color cues for grouping by current models is further highlighted on a grayscale variant of the CLEVR dataset where we see that the RF model struggles to group objects using other features (shape and texture), in contrast to SynCx. Our model shows strong performance compared to more sophisticated synchrony-based baselines [25, 24] on dSprites and CLEVR without the need for additional supervision (‘depth masks’), gating mechanisms or contrastive training. Starting with our model as a simple template, future directions to explore include the use of a temporal difference style loss to weight reconstruction targets and incorporate spatial priors into the binding mechanism. The process of extracting discrete object assignments from continuous phase maps by clustering could be improved by accounting for outliers, non-Gaussian distributed phase values with unequal sizes and variances.
@@ 1 $\times$ 1 conv, 3 channels, stride 1
We seek to visualize the complex-valued feature map $\mathbf{h}\in\mathbb{C}^{h^{\prime}\times w^{\prime}\times d_{\text{out}}}$ at some intermediate layer of the decoder module as an image. We apply dimensionality reduction along the channels of $\mathbf{h}$ so as to represent it as an one dimensional feature map. We start by constructing a complex-valued feature map in polar form with phase components $(\bm{\phi}_{h})$ and magnitude components being the identity, i.e., $\bm{\mu}_{h}\leftarrow\mathbf{1}\in\mathbb{R}^{h^{\prime}\times w^{\prime}\times d_{\text{out}}}$. Then, we convert the feature map obtained in the previous step from polar to Cartesian form. Next, we apply t-SNE to project the Cartesian domain feature map in two dimensions at each spatial location. We use t-SNE implementation in the scikit-learn library [69] (sklearn.manifold.TSNE) with n_iter set to 500, metric set to Euclidean and perplexity set to 15. We recover ‘composite’ phase $\overline{\bm{\phi}}_{h}\in\mathbb{R}^{h^{\prime}\times w^{\prime}}$ by converting the two dimensional feature back to the polar form. The ‘composite’ magnitude $\overline{\bm{\mu}}_{h}\in\mathbb{R}^{h^{\prime}\times w^{\prime}}$ at each spatial location of the map is the Euclidean norm (along channels) of the vector-valued features (i.e., magnitude components of $\mathbf{h}$) at that location. The ‘composite’ phase at every spatial location is visualized as the heatmap value and the ‘composite’ magnitude is its distance from the center in the radial plot. The colors of points are matched across the heatmap and the radial plot such that the color of a pixel in the former correspond to its orientation in the latter. Lastly, the magnitude of background regions are masked out as was the case while extracting object assignments from phase maps.

### v3

@@ 5 Conclusion
We explored synchrony-based binding with an architecture, SynCx, that differs from current models in three respects: i) SynCx’s weights are complex valued, allowing it to encode joint feature-phase configurations in the weights; ii) SynCx is recurrent and stateful, allowing it to perform iterative constraint propagation; and iii) SynCx has an internal representational bottleneck, which require it to use the phases of complex-valued activations to encode statistical regularities in images. These three properties are crucial to extending object-centric phase synchronization behavior to the fully unsupervised setting. Our conceptually elegant model outperforms all state-of-the-art baselines on Tetrominoes and overcomes a common failure mode of these baselines in which similarly colored objects are grouped together. Overreliance on color cues for grouping by current models is further highlighted on a grayscale variant of the CLEVR dataset where we see that the RF model struggles to group objects using other features (shape and texture), in contrast to SynCx. Our model shows strong performance compared to more sophisticated synchrony-based baselines [25, 24] on dSprites and CLEVR without the need for additional supervision such as ‘depth masks’, gating mechanisms or contrastive training. Starting with our model as a simple template, future directions to explore include the use of a temporal difference style loss to weight reconstruction targets and incorporate spatial priors into the binding mechanism. The process of extracting discrete object assignments from continuous phase maps by clustering could be improved by accounting for outliers, non-Gaussian distributed phase values with unequal sizes and variances. While recent synchrony-based models have been steadily closing the gap in grouping performance to well-established slot-based approaches such as SlotAttention [34] on simple synthetic datasets (Tetrominoes and dSprites) the gap remains significant on CLEVR. Further, they are yet to be scaled up to more challenging multi-object visual benchmarks (e.g. MOVi [70]) which offers an open challenge for future work.
@@ 1 $\times$ 1 conv, 3 channels, stride 1
We seek to visualize the complex-valued feature map $\mathbf{h}\in\mathbb{C}^{h^{\prime}\times w^{\prime}\times d_{\text{out}}}$ at some intermediate layer of the decoder module as an image. We apply dimensionality reduction along the channels of $\mathbf{h}$ so as to represent it as an one dimensional feature map. We start by constructing a complex-valued feature map in polar form with phase components $(\bm{\phi}_{h})$ and magnitude components being the identity, i.e., $\bm{\mu}_{h}\leftarrow\mathbf{1}\in\mathbb{R}^{h^{\prime}\times w^{\prime}\times d_{\text{out}}}$. Then, we convert the feature map obtained in the previous step from polar to Cartesian form. Next, we apply t-SNE to project the Cartesian domain feature map in two dimensions at each spatial location. We use the t-SNE implementation in the scikit-learn library [71] (sklearn.manifold.TSNE) with n_iter set to 500, metric set to Euclidean and perplexity set to $20$. We also experimented with UMAP [72] for the dimensionality reduction computation but found it to be inferior to t-SNE qualitatively (see Figure 10) and processing time per image (UMAP takes $\approx$ 24 seconds whereas t-SNE takes $\approx$ 8 seconds). We use the UMAP implementation in the umap-learn library [73] (umap.UMAP) with default arguments. Therefore, we use t-SNE for dimensionality reduction in the phase visualization process. We recover ‘composite’ phase $\overline{\bm{\phi}}_{h}\in\mathbb{R}^{h^{\prime}\times w^{\prime}}$ by converting the two dimensional feature back to the polar form. The ‘composite’ magnitude $\overline{\bm{\mu}}_{h}\in\mathbb{R}^{h^{\prime}\times w^{\prime}}$ at each spatial location of the map is the Euclidean norm (along channels) of the vector-valued features (i.e., magnitude components of $\mathbf{h}$) at that location. The ‘composite’ phase at every spatial location is visualized as the heatmap value and the ‘composite’ magnitude is its distance from the center in the radial plot. The colors of points are matched across the heatmap and the radial plot such that the color of a pixel in the former correspond to its orientation in the latter. Lastly, the magnitude of background regions are masked out as was the case while extracting object assignments from phase maps.
@@ Appendix C Additional Visualizations
Figure 10: Collection of samples comparing the use of t-SNE versus UMAP for the dimensionality reduction computation in the visualization process. Panels 1 and 5 show cases where the clusters of projected phase maps of both t-SNE and UMAP correlate with the number of objects in the image. Panels 2, 3 and 4 show cases where the clusters of phases projected using t-SNE correlate with the number of objects in the image whereas the clusters of phases projected using UMAP do not. Overall, we find that t-SNE produces qualitatively better projections than UMAP, i.e., phase clusters correlate better to the number of objects in the image.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: In the abstract we make the following claims: "recurrent architecture with complex-valued weights perform equally or better at unsupervised object discovery compared to architectures that have complex activations but real-valued weights" which is supported by superior performance in the experiments; we claim that SynCx (our model) avoids certain systematic grouping errors of current models, such as the inability to separate similarly colored objects without additional supervision, which is also supported by experiments with grayscale data.
  5. 2.
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
  8. Justification: We discuss limitations of our method in a dedicated paragraph (last paragraph of the Experiments section), in particular we show its robustness to grouping in the absence of color information.
  9. 3.
    Theory Assumptions and Proofs
  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  11. Answer: [N/A]
  12. 4.
    Experimental Result Reproducibility
  13. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  14. Answer: [Yes]
  15. Justification: We provide all architectural and hyperparameter details in the paper.
  16. 5.
    Open access to data and code
  17. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  18. Answer: [Yes]
  19. Justification: Data we used is already public and upon acceptance we will publish our code.
  20. 6.
    Experimental Setting/Details
  21. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  22. Answer: [Yes]
  23. Justification: All training and test details are provided in Appendix A.
  24. 7.
    Experiment Statistical Significance
  25. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  26. Answer: [Yes]
  27. Justification: All experimental results are provided with mean and standard deviation across 5 seeds.
  28. 8.
    Experiments Compute Resources
  29. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  30. Answer: [Yes]
  31. Justification: We provide sufficient information (see Appendix A) of the hardware resources required to reproduce the main results of the paper as well resources used during the prototyping phase.
  32. 9.
    Code Of Ethics
  33. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  34. Answer: [Yes]
  35. Justification: We have reviewed the NeurIPS Code of Ethics and confirm that our work respects it, i.e. it does not involve human subjects, the data we use is synthetic (no privacy or ethical concerns) and we foresee no harmful consequences of our work as the models are small scale and training and evaluation data is all synthetic.
  36. 10.
    Broader Impacts
  37. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  38. Answer: [N/A]
  39. Justification: We foresee neither negative nor positive societal impacts of our work as the visual scenes we experiment on are highly synthetic (toy tasks) nowhere near the the complexity or scale of naturalistic images/videos. We also only use very small-scale models (less than 1M params) which are not mature yet for use in production.
  40. 11.
    Safeguards
  41. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  42. Answer: [N/A]
  43. Justification: We do not use language models, image generators or datasets scraped from the internet in our work, so we foresee no risk of misuse.
  44. 12.
    Licenses for existing assets
  45. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  46. Answer: [Yes]
  47. Justification: The SynCx model is an original contribution of ours. The baseline models CAE, CAE++, CtCAE and RF have been properly credited with citations to the research papers and code repositories. We also properly cite the data source for the multi-object suite.
  48. 13.
    New Assets
  49. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  50. Answer: [N/A]
  51. Justification: We don’t introduce any new assets.
  52. 14.
    Crowdsourcing and Research with Human Subjects
  53. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  54. Answer: [N/A]
  55. Justification: We do not have any crowd sourcing or research with human subjects.
  56. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  57. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  58. Answer: [N/A]
  59. Justification: We do not have any research with human subjects in our work.